suppose in an experiment data is collected or a variation of current with resistance r is the plot of our business I in this is made how does one determine the value of V from that graph explain your answer is appropriate expression
1
Expert's answer
2021-07-12T02:30:02-0400

Since according to Ohm's law the voltage is


"V=RI,"

just take resistance and current at the required point and multiply them.


Graph






Other graph collection of data




"R=\frac{∆V}{∆I}"

"R=\frac{10.2-3.4}{3-2}=6.8\Omega"